Please solve 2x+5-(3x+2)=4x
Answer by ewatrrr(24785)

 
Hi,
2x+5-(3x+2)=4x
2x+ 5 - 3x - 2 = 4x | Combining Like Terms
3 = 5x
3/5 = x